If a ball is thrown from a height and it falls on the ground, how are we to calculate the time taken? The projectile part till that height can be calculated by 2usina/g but the rest of the motion, ie falling from that height to the ground how do we calculate that?

Grade:11

1 Answers

aditya kulkarni
6 years ago
We can use energy conservation fro this purpose.
Consider that the ball is initially thrown with velocity ‘u’ from a height ‘h’.
It will have its final velocity ‘v’ just before hitting the ground.
So we can use energy conservation for the situation-
therefore      total initial energy = total final energy;
(1/2mu2+mgh)=(1/2mv2      where m= mass of the ball; u=initial velocity; v=final velocity before hitting the ground; h=height from where the ball is thrown; g= gravitational accleration
NOTE:- here we have taken ground as the reference. 
 
 
 
If you agree with my answer please click “Approve” below.